OpenStudy (anonymous):

Simplify by taking roots of the numerator and denominator. sqrt16/81

jimthompson5910 (jim_thompson5910):

Hints: \[\sqrt{\frac{16}{81}}=\frac{\sqrt{16}}{\sqrt{81}}\] \[\sqrt{16}=\sqrt{4^2}=4\] \[\sqrt{81}=\sqrt{9^2}=9\]